George Soros and Warren Buffet both believe the stock market is going down (i.e. stock prices will fall). Buffet decides to use an option strategy and Soros decides to use a forwards strategy. Currently, the stock market is priced at $200. A call option has a strike of $205 and costs $2. A put option has a strike of $195 and costs $3. The forward price is $201. Each investor uses the respective derivative strategy mentioned above. What is the profit/loss of each investor if the market finishes at $198?

Answers:

  • Soros profits $2

    Buffet profits $3

  • Soros breaks even

    Buffet profits $1

  • Soros loses $2

    Buffet profits $1

  • Buffet loses $3

    Soros profits $3

  • Soros profits $3

    Buffet loses $2
Answer #1

Both believes that the market will go down .

Soros will sell a forward at forward price of 201. If the stock price falls below 201 Soros will profit else he will lose if the stock price goes above the forward price

Warren will purchase a put option at strike rate of $195 and has to pay premium of $3

Stock price goes to $198

Profit for Soros = Forward price - stock price = 201- 198 = $3

Profit for warren = MAX ( Strike price - stock price, 0) - option premium

= MAX ( 0,0 ) - 3

= - 3

Warren will let the option expire as it is out of money.

So Soros profits $3 and warren loses $3